The championship record in baseball in a certain year was 52 games won, and 14 lost. What percent of the games were won?
Answer by greenestamps(13200)



52 wins; a total of 52+14=66 games; so they won 52 out of 66:
52/66 = 26/33 = 78/99 = .78787878... = 78.8% to 1 decimal place.